# Towards Fully FP8 GEMM LLM Training at Scale

### v1

@@ 5.2 Downstream performance
We compare our proposals with the higher-precision Llama3 baseline across a wide range of standard benchmarks to measure their downstream performance. In Table 3, we report some of the most relevant scores along with an average* across a larger set of tasks, detailed in the Appendix. All FOG variants offer comparable downstream performance with the higher precision Llama3 baseline with FOG-max architecture, even outperforming it. The 1.5B models are trained on 125B tokens, whereas smaller models are trained on 50B tokens. We note that we initially tested the idea of cooling down the previously constant weight decay during the learning cooldown phase, aiming to conserve model weights’ norm [15]. We show later in the Appendix that such intervention has no noticeable effect on stability neither performance. We had to conserve it for consistency and fair comparison across ablations. Finally, we use identical hyper-parameters such as batch size, learning rate and weight decay.
@@
Table 3: Performance across various tasks. For each task and model size, the first score results from the BF16 ablation and the second from the FP8DPA one.
5.3 Efficiency
Table 4 explores the efficiency of our architectures in the 1.5B and 8B model scales for FP8 training. We compare our FP8DPA solution with the BF16 baseline and the FP8 training enabled by the use of SmoothSwiGLU, which is, to the best of our knowledge, the only dense architecture proposal demonstrated to work at scale with FP8. Note that the SmoothSwiGLU cannot benefit from enabling FP8 GEMMs in the attention mechanism, as it was shown to suffer a big loss divergence in Figure 6. Furthermore, our throughput gains increase as the model size increases. The GEMM input tensors increase in size and consume significantly more time during the overall forward-backward pass, compared with other operators.
Size Model Precision
Throughput
(tokens/second/GPU)
8B Llama BF16 9105
FOG-max FP8DPA 12344 (+35.5%)
FOG-opt FP8DPA 12414 (+36.3%)
FOG-flash FP8DPA 12764 (+40.2%)
Llama+SmoothSwiGLU FP8 12228 (+34.3%)
1.5B Llama BF16 46470
FOG-max FP8DPA 53551 (+15.2%)
FOG-opt FP8DPA 53877 (+15.9%)
FOG-flash FP8DPA 54848 (+18.0%)
Llama+SmoothSwiGLU FP8 54903 (+18.1%)
Table 4: Throughput measures with FOG versus other baselines. Using eight 4xGH200 nodes with Zero-1 sharding [24], without model parallelism, for 8B models and a single 4xGH200 node for 1.5B models. Notably, in the 8B scale, all FOG variants outperform other architectures.
5.4 Long-data regimes and FP16 optimizer states
To further justify the viability of FP8DPA long training with FOG, we train our 1.5B FOG-max to 420B tokens, way beyond the previously identified 200B tokens divergence mark of Llama2-7B [6]. Note that our observation of smaller models’ tendancy to diverge earlier with FP8DPA, and the long-term outlier analysis in Section 4 further proves the sufficiency of such a training duration. We also switch to use FP16 optimizer states and BF16 gradients after 130B tokens, saving up memory previously used by full precision states, gradients, and model parameters master copy. We display the learning dynamics of our approach in Figure 7. The language modeling loss exhibits equal to better smoothness compared to the corresponding Llama baseline.
Figure 7: Long-data training regimes. FOG-max 1.5B FP8DPA is trained on 420B tokens. The higher precision Llama3 shorter experiment is also included as reference.
@@ 6 Limitations
Despite its robustness, outstanding throughput and relative flexibility, the FOG set of architectures remains mainly limited by two aspects. Since the introduction of transformers [30], various iterations have followed for more stability and higher performance. Future better variants might be incompatible with our design and thus require different recipes to ensure FP8DPA training. Second and more importantly, FOG is trained with BF16 GEMM for the final projection. This operator is known to be very sensitive to outliers and has always been used with half-precision in forward-backward FP8 training approaches, including ours.
Moreover, this work focuses on maximising throughput gains while ensuring both strong stability and downstream performance. We also show it to be robust to half-precision optimizer states and gradients, allowing considerable memory savings. Yet, we do not discuss further casting down optimizer states to FP8. Although this will allow for comparatively less memory savings, especially as the optimizer states are usually sharded [24], it may help to avoid some memory saturation issues in compute-constrained settings. Finally, even though we managed to scale training to a long data regime at 1.5B scale, showed consistent observations across model sizes, and provided further guarantees, our experiments were limited to a short regime of 20B tokens at the 8B parameter scale. Due to computational constraints, we decided to keep the study of these limitations for future work.
@@ 7 Conclusion
In this paper we demonstrate, for the first time, stable LLM training with fully FP8 computations within the transformer blocks–including the attention mechanism–without sacrificing performance. We tested FP8DPA training across a wide set of previously proposed architectures and show that they consistently diverge early during training, highlighting the difficulty of FP8DPA training and novelty in our results. Moreover, in contrast with other granular scaling recipes, we use the low-overhead delayed scaling FP8 strategy. Our design provides on-par downstream quality with the higher precision baseline, while offering up to 40% faster training. We scale our 1.5B model to 420B tokens, 14x the Chinchilla-optimal data budget for its size. Our work brings the community one step closer to fully FP8 GEMM training at scale i.e including the language modeling head.
We further justify the long-term stability of our architecture by observing the outlier training dynamics across key activations by using kurtosis. The use of kurtosis to track outliers present during training was shown to provide meaningful insights to favour certain architectural components or to predict future instabilities, as it is a quantitative metric that measures outliers.
@@ Appendix A Hyperparameters
We detail the selection of hyperparameters used in Table 5. For the case of FOG-flash, the $\alpha_{0}$ initialization value of $\tanh_{\alpha}$ entropy-regularization is 0.5 for all model sizes. All models use a linear warmup schedule, and 1-sqrt cooldown schedule. The long-data 1.5B FOG-max experiment was trained for a total of 400,000 steps, consuming approximately 419.4B tokens, using the same hyperparameters as the shorter run–except for the absence of a cooldown phase.
@@
Table 5: Hyperparameters used in experiments. Note that FFN hidden size indicates the dimensionality of each linear projection in gated activation functions; networks without GLUs use 1.5×1.5$\times$1.5 × this value to match the parameter count. Softmax scale specified only applies to FOG models, all other models follow the standard $s=1/\sqrt{D_{QK}}$.
@@ Appendix B Architectures
    An input scaling factor $u\in(0,\infty)$, which may equal $1$
@@
OP${}^{(\text{a})}$ $\sigma_{0}^{-1}$ $\operatorname{id}$ $\operatorname{id}$ $\mathrm{LayerScale}_{\boldsymbol{\gamma}}$ $\mathcal{N}_{\boldsymbol{\gamma}}$ $\mathrm{GeLU}$
FOG-max${}^{(\text{a,b})}$ $\sigma_{0}^{-1}$ $\operatorname{id}$ $\operatorname{id}$ $\mathcal{N}_{\boldsymbol{\gamma}}$ $\mathcal{N}$ $\mathrm{xIELU}$
FOG-opt${}^{(\text{a,b})}$ $\sigma_{0}^{-1}$ $\operatorname{id}$ $\operatorname{id}$ $\mathcal{N}_{\boldsymbol{\gamma}}$ $\mathcal{N}$ $\mathrm{GeLU}$
FOG-flash${}^{(\text{a,b})}$ $\sigma_{0}^{-1}$ $\operatorname{id}$ $\operatorname{id}$ $\mathcal{N}_{\boldsymbol{\gamma}}$ $\tanh_{\alpha}$ $\mathrm{GeLU}$
@@
Appendix C FP8 training
For all our experiments, we used Transformer Engine’s delayed scaling implementation with history length $\ell=1024$ and margin $m=0$. Mathematically, given a history of abs-max values, denoted $H=\{h_{t}\}_{t=1}^{\ell}\subseteq[0,\infty)$, of a tensor $\mathbf{X}$, we define its scaling factor as:
@@ Appendix C FP8 training
[…]
@@
[…]
@@
Table 7: Comparison between FP8 methods. The FP8DPA method allows for all—excluding the output head—GEMMs computations to be done with FP8 precision. In contrast, FP8 training uses higher precision for all GEMMs in the core attention. The linear operators are linear layers of the form $\mathrm{Linear}_{\mathbf{W}}(\mathbf{X})=\mathbf{XW}$: namely the FFN linear layers, QKV projections and attention output projection. See Equation (1) for the definition of the attention probability matrix $\mathbf{P}$.
Appendix D Evaluations
We selected the following set of benchmarks: ARC-Easy, CommonsenseQA, HellaSwag, LAMBADA-OpenAI, LAMBADA-standard, OpenBookQA, PIQA, SocialIQA, and WinoGrande.
We used a standard open-source LLM evaluation package for conducting these evaluations, as cited in the code repository https://github.com/anonymous4375934/FOG.
In Table 3, we report raw accuracy scores as percentages on three key benchmarks as well as the average over the full set of tasks mentioned above. In the following table 8, we provide all scores along with their estimation errors for the 1.5B model size, demonstrating that the slight differences observed across many values are statistically insignificant.
@@ Appendix D Evaluations
PIQA $72.5\mid-$ $72.6\mid 73.3$ $72.5\mid 72.0$ $72.2\mid 72.0$$\pm 1.0$
Commonsense-qa $19.6\mid-$ $20.2\mid 22.2$ $19.3\mid 21.2$ $21.1\mid 20.8$$\pm 1.2$
Lambada-openai $44.5\mid-$ $43.7\mid 44.6$ $44.3\mid 44.5$ $42.4\mid 41.2$$\pm 0.7$
Lambada-standard $38.9\mid-$ $37.0\mid 39.5$ $37.7\mid 37.9$ $35.9\mid 33.8$$\pm 0.7$
@@
Weight decay cooldown As mentioned in Section 5.2, we experimented with cooling down the weight decay, often used as a constant value equal to $0.1$ that is coupled with the learning rate, to see if it solves the OP+frozenQK architecture’s consistent divergence during the learning rate decay phase. We also tested it on other architectures and, to optimize the use of resources, we had to keep it later for the final experiments. This trick helped stabilize the weights’ norm indeed, but couldn’t solve the divergence issue. However, it had no effect on final performance nor on stability. The following table highlights this no-effect claim at 1.5B scale.
@@
FOG-opt cooldown converges 46.0 ±$\pm 0.3$
FOG-opt constant converges 46.3 ±$\pm 0.3$
@@
Appendix E Computational Resources
Our experiments were conducted on nodes equipped with 4 Grace Hopper GPUs each. We typically used 4, 8, and 16 nodes for our 390M, 1.5B, and 8B parameter experiments, respectively, with minor variations across different runs. Importantly, all throughput measurements were taken under identical hardware configurations. Table 10 details the computational resources in GPU hours (GPUh) required for our main experimental results. This includes the computational cost of training all architectures that diverged during FP8DPA training, the FP8DPA and BF16 stable training runs for our three architectures, and the BF16 Llama3 baseline. The aggregation includes node start-up times, computation lost due to node failures, and overhead from calculating and logging kurtosis metrics. The complete research project required additional computational resources beyond those specified in the table, as we conducted numerous preliminary experiments and explored ideas that did not appear in the final paper.

### v2

@@ 5.3 Downstream performance
We compare our proposals with the higher-precision Llama3 baseline across a wide range of standard benchmarks to measure their downstream performance. Inference during down-stream evaluation uses BF16 precision. In Table 5, we report some of the most relevant scores along with an average across a larger set of tasks, detailed in the Appendix G. All FOG variants offer comparable downstream performance with the higher precision Llama3 baseline with FOG-max architecture, even outperforming it. The 1.5B models are trained on 125B tokens, whereas smaller models are trained on 50B tokens.
@@
Table 5: Performance across various tasks. For each task and model size, the first score results from the BF16 ablation and the second from the FP8DPA one. The average* is across a larger set of tasks, show in Appendix G.
5.4 Long-data regimes
To further justify the viability of FP8DPA long training with FOG, we train a 1.5B FOG-max on 450B tokens, way beyond the previously identified 200B tokens divergence mark of Llama2-7B [6]. Note that our observation of smaller models’ tendancy to diverge earlier with FP8DPA, and the long-term outlier analysis in Section 4 further underline the sufficiency of such a training duration.
We also switch to use FP16 optimizer states and BF16 gradients after 130B tokens, saving up memory previously used by full precision states, gradients, and model parameters master copy. We display the learning dynamics of our approach in Figure 7. The language modeling loss exhibits equivalent to better smoothness compared to the corresponding Llama baseline.
Figure 7: Long-data training regimes. FOG-max 1.5B FP8DPA is trained on 450B tokens. The higher precision Llama3 experiment is included as reference. Note that a learning rate cooldown is performed during the last 25B tokens of each experiment, following the WSD schedule. Loss reported is the 200-rolling-window mean and 5%-95% quantile bands.
5.5 Additional results
SwiGLU
This work mainly studies three architectural variants, sharing in common the use of point-wise activation functions. As mentioned in Section 4, we extended FOG to gated activation functions by showing its stability when using a gated MLP (SwiGLU). Details in Appendix D.
FP8 optimizer moments
We ran an additional FP8DPA experiment using FOG-flash following the same 390M scale setup but with 8 bit optimizer moments. The loss converged smoothly to a value of $2.645$, nearly identical to the value $2.649$ obtained with higher-precision moments. More details are available in Appendix E.
MoE
We tested FOG under a Mixture-of-Experts (MoE) setting and the training was consistently stable. This experiment further supports the robustness of FOG. More details can be found in Appendix D.
@@ 6 Limitations
Despite its robustness, record throughput boosts, and flexibility, the FOG set of architectures remains bounded by the following limitation. The final projection (LM head) is still performed in BF16. This operator is known to be very sensitive to outliers and has been used with half-precision in forward-backward FP8 training approaches, including ours. Due to computational constraints, we decided to keep the study of this limitation for future work.
@@ 7 Conclusion
In this paper we demonstrate, for the first time, stable LLM training with fully FP8 matrix multiplications within the transformer blocks–including the attention mechanism–without sacrificing performance. We tested FP8DPA training across a wide set of previously proposed architectures and show that they consistently diverge early during training, highlighting the difficulty of FP8DPA training and novelty in our results. Moreover, in contrast with other granular scaling recipes, we use the low-overhead delayed scaling FP8 strategy. Our design provides on-par downstream quality with the higher precision baseline, while offering up to 43% faster training at 8B scale. We scale our 1.5B model to 450B tokens, 15x the Chinchilla-optimal data budget for its size. Our work brings the community one step closer to fully FP8 GEMM training at scale i.e including the language modeling head. We further justify the long-term stability of our architecture by observing the outlier training dynamics across key activations by using kurtosis. The use of kurtosis to track outliers present during training was shown to provide meaningful insights to favour certain architectural components or to predict future instabilities, as it is a quantitative metric that measures outliers.
@@ Appendix A Hyperparameters
We detail the selection of hyperparameters used in Table 6. For the case of FOG-flash, the $\alpha_{0}$ initialization value of $\tanh_{\alpha}$ entropy-regularization is 0.5 for all model sizes. All models use a linear warmup schedule, and 1-sqrt cooldown schedule. The long-data 1.5B FOG-max experiment was trained for a total of 430,000 steps, consuming approximately 450.9B tokens, using the same hyperparameters as the shorter run, including warmup steps.
@@
Table 6: Hyperparameters used in experiments. Note that FFN hidden size indicates the dimensionality of each linear projection in gated activation functions; networks without GLUs use $1.5\times$ this value to match the parameter count. Softmax scale specified only applies to FOG models, all other models follow the standard $s=1/\sqrt{D_{QK}}$.
@@ Appendix B Architectures
    An input scaling factor $u\in(0,\infty)$, which may equal 11
@@
[…]
@@
xIELU activation function
Introduced in [13], the xIELU activation function is defined element-wise as:
$\mathrm{xIELU}(x):=\begin{cases}\alpha_{p}x^{2}+0.5x&\text{if $x>0$,}\\ \alpha_{n}(e^{x}-1)-\alpha_{n}x+0.5x&\text{if $x\leq 0$.}\end{cases}$ (2)
where $\alpha_{p}$ and $\alpha_{n}$ are trainable scalars per layer. xIELU is an extension of Squared ReLU and has been adopted and validated at scale [11].
Appendix C Outliers impact quantization
We begin with a useful definition.
$\tau$-outlier: Given $\mathbf{x}\in\mathbb{R}^{d}$ and $\sigma=\text{rms}(\mathbf{x})$, the element $x$ of $\mathbf{x}$ is a $\tau$-outlier if $|x|\geq\tau\sigma$.
As $\tau$ increases, $x$ becomes a larger outlier ($\sigma$ represents the natural magnitude of $\mathbf{x}$). In practice, $\tau\gg 1$. Before FP8 quantization, each tensor is scaled with $s(\mathbf{x}):=\text{MaxFP8Value}/\text{absmax}(\mathbf{x})$ to better utilize limited FP8 dynamic range.
Theorem 1.
Let $\mathbf{x}\in\mathbb{R}^{d}$ have $\tau$-outlier $x_{j}$, and $\mathbf{x}^{\prime}\in\mathbb{R}^{d}$ have a $\tau^{\prime}$-outlier $x^{\prime}_{j}$ with $\tau^{\prime}>\tau$. Then for any subset $T\subseteq\{1\ldots d\}\setminus j$, vector $\mathbf{x}^{\prime}_{T}$ will be quantized less accurately than $\mathbf{x}_{T}$.
In other words, larger outlier values lead to less precise FP8-quantized results.
C.1 Proof
Let $r=\text{FP8MaxValue}$, $s=r/\text{absmax}(\mathbf{x})$, and $s^{\prime}=s(\mathbf{x}^{\prime})$.
Since $\text{absmax}(\mathbf{x})\geq|x_{j}|\geq\tau\sigma$, then $s\leq\frac{r}{\tau\sigma}$ (similarly, $s^{\prime}\leq\frac{r}{\tau^{\prime}\sigma}$). Let $m=\text{absmax}(\mathbf{x}_{T})$ and $m^{\prime}=\text{absmax}(\mathbf{x}^{\prime}_{T})$.
Elements of $\mathbf{x}_{T}$ lie in [m,m][-m,m]. After scaling, the range becomes $rm/(\tau\sigma),rm/(\tau\sigma)]$ in $s\mathbf{x}_{T}$, and $rm/(\tau^{\prime}\sigma),rm/(\tau^{\prime}\sigma)]$ in $s^{\prime}\mathbf{x}^{\prime}_{T}$. Since $\tau^{\prime}>\tau$, $rm/(\tau^{\prime}\sigma),rm/(\tau^{\prime}\sigma)]\subsetneq[-rm/(\tau\sigma),rm/(\tau\sigma)]$, so $\mathbf{x}^{\prime}_{T}$ has smaller range.
This narrowed range contains fewer n-bit representable numbers, proving the theorem.
The proof applies to any subset $T$, including the set of “typical” values (e.g., 90%-quantile).
Theorem 1 guarantees large outliers worsen quantization on 90% of tensor elements.
C.2 Empirical confirmation
We measured activation values of Llama and FOG-max 1.5B during mid-training on a micro batch of data (precisely the second FFN layer’s input, before quantization).
Observation: Llama presents a 688-outlier while FOG-max shows only a 183-outlier. Using 90%-quantile, we get that 90% of Llama’s activation coefficients scale to [0.289,0.289][-0.289,0.289] range, while FOG-max allows a much broader range of [2.084,2.984][-2.084,2.984].
Appendix D FOG extensions
FOG-SwiGLU
In addition to our main experiments, we trained a 1.5B FOG model using the SwiGLU activation function, which we label FOG-SwiGLU. This architecture was adapted from FOG-max, changing the activation function to SwiGLU and adjusting FFN hidden size to match the parameter count. Figure 8 shows the loss progression of this model under FP8DPA training. This experiment resulted in a completely stable training, and further demonstrates the flexibility of activation functions suitable in our design.
Figure 8: FOG-SwiGLU 1.5B FP8DPA run. FOG-max included as reference. We observe stable training dynamics for both approaches. The reported loss is the 200-rolling-window mean and 5%-95% quantile bands.
MoE extension
We adapted FOG-flash architecture to follow an MoE design, keeping the backbone configuration of the 390M model (hidden size, number of layers, etc), but upscaling with 8 FFN experts (2 active) following [17], resulting in a 1.8B model, trained from scratch under the same configuration as all 390M models. Additionally, we employ the z-loss [17] with coefficient of 0.01, and no explicit loss-balancing loss, but rather an expert bias [33] with update rate of 0.01. We trained this model under BF16 and FP8DPA training, resulting on a final converged loss of 2.477 and 2.483, respectively, as shown in Figure 9. This stable training result suggests good generalization capabilities for FOG to other MoE designs under FP8DPA training. We further adapted our 1.5B models and similarly scale up to 8 experts (2 active) to measure throughput gains at a larger scale. Table 8 summarizes our results, where FOG still provides the most throughput gains.
Figure 9: FOG-flash-MoE 1.8B-8E FP8DPA loss progress. Both BF16 and FP8DPA trainings are shown. The FP8DPA training remains stable for the entire duration of training, and the final loss converged to in both precisions remains within $\pm 0.005$, suggesting comparable downstream capabilities. Loss reported is the 200-rolling-window mean and 5%-95% quantile bands.
Model Precision
Throughput
(tokens/second/GPU)
Llama3 BF16 3336
Llama+SmoothSwiGLU FP8 4202 (+24.8%)
FOG-flash FP8DPA 4351 (+29.2%)
Table 8: MoE training throughput. Measurements of 41B-8E MoEs taken using 4xGH200 nodes with expert parallel and pipeline parallel size of 4 using a batch size of 512. As with dense models, FOG-flash outperform all other architectures.
Appendix E FP8 training
In all our experiments, we used Transformer Engine’s delayed scaling implementation with history length $\ell=1024$ and margin $m=0$. Mathematically, given a history of abs-max values, denoted $H=\{h_{t}\}_{t=1}^{\ell}\subseteq[0,\infty)$, of a tensor $\mathbf{X}$, we define its scaling factor as:
@@ Appendix E FP8 training
where $\mathrm{FP8MaxValue}\in(0,\infty)$ is the maximum value representable with the FP8 format used. We update the history using $H\leftarrow\{\max_{x\in\mathbf{X}}|x|\}\cup\{h_{t}\}_{t=2}^{\ell}$ to use for this activation in the next iteration. The end-to-end FP8 matrix multiplication is
@@
where $\mathrm{FP8GEMM}$ receives FP8 tensors and returns the BF16 result. We further detail the precision used for every matrix multiplication during our FP8 and FP8DPA experiments in Table 9.
@@
Table 9: Comparison between FP8 methods. The FP8DPA method allows for all GEMM computations—excluding the output head— to be done with FP8 precision. In contrast, FP8 training uses higher precision for the core attention computation. The linear operators are linear layers of the form $\mathrm{Linear}_{\mathbf{W}}(\mathbf{X})=\mathbf{XW}$: namely the FFN linear layers, QKV projections and attention output projection. See Equation (1) for the definition of the attention probability matrix $\mathbf{P}$.
FP8 optimizer moments
To further reduce memory usage, we tested FOG-flash under the usual FP8DPA setting, with an additional constraint: FP8 optimizer moments. This extends the typical setting of half-precision gradients and moments used in most of our experiments. Figure 10 shows the training loss across three different settings for comparison.
Figure 10: FOG-flash 390M loss curve comparing training precision. Our design reaches similar loss when trained with either precision. Loss reported is the 200-rolling-window mean.
Fine-grained scaling recipes
Recent FP8 training achievements, such as DeepSeek’s DeemGEMM kernels, involve the use of fine-grained FP8 scaling recipes to provide a more robust training regime. While these options could potentially enable FP8 training when tensor-wise scaling alternatives diverge, it comes with a significant overhead. We validate this claim using TransformerEngine’s Blockwise scaling at the 8B scale. Training throughput is reported in Table 10. Using FP8DPA training with delayed scaling recipe provides the highest boost across all tested methods.
Model Precision FP8 Recipe
Throughput
(tokens/second/GPU)
Llama3 BF16 N/A 9.48k
Llama3 FP8 Blockwise 11.18k (+17.9%)
OP FP8 Delayed 12.14k (+28.1%)
Llama+SmoothSwiGLU FP8 Delayed 13.1k (+38.2%)
FOG-flash FP8DPA Delayed 13.52k (+42.6%)
Table 10: Training throughput. Measurements taken using eight 4xGH200 nodes with Zero-1 sharding [26], without model parallelism using a batch size of 1024. Notably FOG-flash outperform all other architectures.
Appendix F Long context
Note that when global batch size (GBS) increases –micro batch size fixed–, computation time takes over communication time. Therefore, the larger the GBS, the higher the throughput gains for all approaches: FOG-flash reaches +42.6% with GBS=1k compared to +40.2% with GBS=512. Moreover, enabling FP8 computations in the attention bring unique throughput benefits under long-context training. Hence the large efficiency gap achieved by FOG-flash FP8DPA compared to Llama3+SmoothSwiGLU in Table 11.
Context Length TP Llama3 (BF16) FOG-flash (FP8DPA) Llama3+SmoothSwiGLU (FP8)
4096 1 9.48K 13.52K 13.13K
8192 1 9.08K 13.03K Out Of Memory
8192 2 7.49K 10.42K 10.05K
16384 2 6.81K 9.45K 8.93K
Table 11: Training throughput under varying sequence lengths, performed at 8B scale using eight 4xGH200 nodes with a global batch size of 1024. TP refers to Tensor Parallelism.
Appendix G Evaluations
We selected the following set of benchmarks: ARC-Easy, CommonsenseQA, HellaSwag, LAMBADA-OpenAI, LAMBADA-standard, OpenBookQA, PIQA, SocialIQA, and WinoGrande. We used a standard open-source LLM evaluation package for conducting these evaluations, as cited in the code repository https://github.com/anonymous4375934/FOG.
In Table 5, we report raw accuracy scores as percentages on three key benchmarks as well as the average over the full set of tasks mentioned above. In Table 12, we provide all scores along with their estimation errors for the 1.5B model size, demonstrating that the slight differences observed across many values are statistically insignificant.
@@ Appendix G Evaluations
PIQA $72.5\mid$- $72.6\mid 73.3$ $5\mid 72.0 72.27$$2\mid 72.0$ $\pm 1.0$
Commonsense-qa $19.6\mid$- $20.2\mid 22.2$ $19.3\mid 21.2$ $1\mid 20.8 ±1.2\pm 1.2$
Lambada-openai $44.5\mid$- $3.7\mid 44.6 44$$4.3\mid 44.5$ $42.4\mid 41.2$ $\pm 0.7$
Lambada-standard $38.9\mid$- $0\mid 39.5 37.73$$7\mid 37.9$ $35.9\mid 33.8$ $\pm 0.7$
@@
Weight decay cooldown As mentioned in Section 3, we experimented with cooling down the weight decay, often used as a constant value equal to $0.1$ that is coupled with the learning rate, to see if it solves the OP+frozenQK architecture’s consistent divergence during the learning rate decay phase. We also tested it on other architectures and, to optimize the use of resources, we had to keep it later for the final experiments. This trick helped stabilize the weights’ norm indeed, but couldn’t solve the divergence issue. Moreover, it had no effect on final performance nor on stability. Table 13 highlights this no-effect claim at 1.5B scale.
@@
FOG-opt cooldown converges 46.0 $\pm 0.3$
FOG-opt constant converges 46.3 $\pm 0.3$
@@
Appendix H Computational Resources
Our experiments were conducted on nodes equipped with 4 Grace Hopper (GH200) GPUs each. We typically used 4, 8, and 16 nodes for our 390M, 1.5B, and 8B parameter experiments, respectively, with minor variations across different runs. Importantly, all throughput measurements were taken under identical hardware configurations. Table 14 details the computational resources in GPU hours (GPUh) required for our main experimental results. This includes the computational cost of training all architectures that diverged during FP8DPA training, the FP8DPA and BF16 stable training runs for our three main architectures, and the BF16 Llama3 baseline. The aggregation includes node start-up times, computation lost due to node failures, and overhead from calculating and logging kurtosis metrics. The complete research project required additional computational resources beyond those specified in the table, as we conducted numerous preliminary experiments and explored ideas that did not appear in the final paper.